OpenStudy (anonymous):

1 question, see below, will give medal.

OpenStudy (anonymous):

@Dobby1

OpenStudy (anonymous):

is it 2,3?

OpenStudy (dobby1):

the solution is the point where the two lines intersect and yes

OpenStudy (anonymous):

note down the point of intersection.

OpenStudy (anonymous):

ok, i thought so haha.

OpenStudy (dobby1):

ok lol great job

OpenStudy (dobby1):

medal?

OpenStudy (dobby1):

thx

OpenStudy (anonymous):

Thanks to you both.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@Dobby1

OpenStudy (dobby1):

hey let me check.

OpenStudy (anonymous):

2.50x+23=2.25x+24 2.50x-2.25x=24-23 .05 x=1 \[\frac{ 5 }{100}x=1 ,calculate x\]

OpenStudy (anonymous):

correction .25 x=1 \[\frac{ 25 }{ 100 }x=1\] calculate x

OpenStudy (dobby1):

oh I hate these what are the other answers?

OpenStudy (dobby1):

b because both equations have either less than or less than or equal to signs meaning both lines have to have the shaded region underneath them neither can have on top and they want the shaded region that both of their shaded regions are inside of.

OpenStudy (anonymous):

so the graph thats half colored in?

OpenStudy (dobby1):

ok let me look I am sorry it takes me a bit to respond I am helping another kid also

OpenStudy (anonymous):

its fine.

OpenStudy (dobby1):

hm I really hate these

OpenStudy (dobby1):

I really suck at explaining these so I am just going to give you the answer

OpenStudy (dobby1):

-14

OpenStudy (anonymous):

Alright, im gonna talk to my teacher and see if she can help me with these because i suck most at these haha

OpenStudy (dobby1):

ok good idea I wish I could explain these but it is really hard to explain them

OpenStudy (dobby1):

b because it is the only one where they have 2 times the amount of gummy candy as jelly beans

OpenStudy (anonymous):

ok.

OpenStudy (dobby1):

no but I can help you with it

OpenStudy (anonymous):

ok.

OpenStudy (dobby1):

4x-7y=-12 for the equation for the top part and -14x+4y=4 for the bottom part so your answer is 4x-7y=-12 -14x+4y=4

OpenStudy (dobby1):

d

OpenStudy (dobby1):

ok \[\frac{ 3x+3y=10 }{ 9x+9y=30 }\]

OpenStudy (dobby1):

that is the first step

OpenStudy (anonymous):

ok

OpenStudy (dobby1):

step 2 12x+12y=40

OpenStudy (dobby1):

step 3 12y=-12x+40

OpenStudy (anonymous):

ok

OpenStudy (dobby1):

step 4 \[y=-x+3\frac{ 1 }{ 3 }\]

OpenStudy (dobby1):

step 5 \[3x+3(-x+3\frac{ 1 }{ 3 })=10\]

OpenStudy (dobby1):

step 6 3x-3x+10=10

OpenStudy (dobby1):

step 7 0+10=10

OpenStudy (dobby1):

final step 10=10

OpenStudy (dobby1):

no sorry

OpenStudy (anonymous):

hello?

OpenStudy (dobby1):

sorry eating lunch

OpenStudy (anonymous):

ok, is that 13 for you?

OpenStudy (dobby1):

no

OpenStudy (anonymous):

can you help me with it?

OpenStudy (dobby1):

i can try

OpenStudy (dobby1):

ok well what they are asking is which one of the points they show you in the graph give P its highest value

OpenStudy (anonymous):

9,0?

OpenStudy (dobby1):

yes because 3(9)+0=27 0+2(8)=16 5(3)+2(4)=23

OpenStudy (anonymous):

so 9,0 is the answer?

OpenStudy (dobby1):

yes

OpenStudy (dobby1):

hm no wow that is so strange

OpenStudy (anonymous):

can you help with this one or too hard?

OpenStudy (dobby1):

i can try

OpenStudy (anonymous):

ok

OpenStudy (dobby1):

oh man I really hate these

OpenStudy (anonymous):

ok well you dont have to do this one, you can help with 12 instead?

OpenStudy (dobby1):

ok thx

OpenStudy (dobby1):

ok do you know how to graph lines.

OpenStudy (anonymous):

no.

OpenStudy (dobby1):

ok it is fairly easy do you have graph paper near you

OpenStudy (anonymous):

no.

OpenStudy (dobby1):

do you have any graph paper at all

OpenStudy (anonymous):

No.

OpenStudy (dobby1):

oh ok you need graph paper to complete this uhm i dont have any either do you have a note book

OpenStudy (anonymous):

alright, nevermind. haha.

OpenStudy (dobby1):

sorry you are going to need to go out and bye some graph paper get some graph paper to finish this test unfortunatly

OpenStudy (anonymous):

Im going out in a bit so il have to pick up some then.

OpenStudy (dobby1):

ok when you get back just ask me to help and I will teach you how

OpenStudy (anonymous):

ok thanks.

OpenStudy (dobby1):

np